# Perturbative soft graviton theorems in de Sitter spacetime

### v2

@@ 6 Soft Graviton Theorems
We must note that our analysis is valid for $\omega l>1$ in the limits of $l\to\infty$ and $\omega\to 0$. To put the soft factor in a proper hierarchy of leading, subleading, and sub-subleading corrections to the flat space soft factor, one can introduce a small parameter $\delta=1/\omega l$. Thus, the soft factor can be expanded in $\omega$ and $\delta$ as follows:
$A^{(0)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l}A^{(1)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l^{2}}A^{(2)}(\{p_{i}\},\omega\hat{k})=\sum_{r=0}^{2}\sum_{s\geq-1}\delta^{r}\omega^{s}S^{(r,s)}(\{p_{i}\},\hat{k}),$ (56)
where $S^{(0,-1)}$ corresponds to the flat space soft factor, and rest of the coefficients are leading, subleading, sub-subleading corrections, and so on.
@@ 6.1 Leading corrections
Let us substitute the scalar field propagator $D(z^{\prime},z)$ into eq. (54) and simplify the expression by keeping the terms up to $\mathcal{O}(l^{-1})$.
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}2\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}\left(1+3\dfrac{p_{i}\cdot z^{\prime}}{2ml}-i\dfrac{(p_{i}\cdot z^{\prime})^{2}}{2ml}-i\dfrac{mz^{\prime 2}}{2l}\right)\\ \cdot e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (57)
To obtain the soft factor in the momentum space, we replace all $z^{\prime\mu}$ by $-ie^{-ip\cdot z^{\prime}}\frac{\partial}{\partial p_{\mu}}e^{ip\cdot z^{\prime}}$ in the above equation.
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}2\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}\\ \cdot\Big{(}1-3i\dfrac{p_{i}^{\mu}}{2ml}\dfrac{\partial}{\partial p^{\mu}}+i\dfrac{p_{i}^{\mu}p_{i}^{\nu}}{2ml}\dfrac{\partial}{\partial p^{\mu}}\dfrac{\partial}{\partial p^{\nu}}-i\dfrac{m}{2l}\dfrac{\partial}{\partial p^{\mu}}\dfrac{\partial}{\partial p_{\mu}}\Big{)}e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (58)
The above expression contains one $\mathcal{O}(l^{0})$ and three $\mathcal{O}(l^{-1})$ terms. We analyze them step by step. Let us first evaluate the $\mathcal{O}(l^{0})$ term:
$\Gamma^{(0)}_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}2\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (59)
By performing the $z^{\prime}$-integral in the above expression, one can generate the delta function $\delta^{(4)}(p-p_{i}-k)$, which can be integrated w.r.t. $p$ to replace all $p$ by $p_{i}+k$ in the expression above. Finally, we take the soft limit, and obtain
$\Gamma^{(0)}_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\int d^{4}ze^{-ip_{i}\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (60)
The momentum $p_{i}$ in the exponential is on-shell, and hence it will remain the same for all the scalars. Thus, we can move $z$-integral outside the summation:
$\therefore\Gamma^{(0)}_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\Gamma^{(0)}_{n}{(\{p_{i}\})},$ (61)
which is the flat space soft graviton theorem. The soft factor is
$A^{(0)}=\sum_{i=1}^{n}\kappa_{i}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}.$ (62)
One can apply the Ward identity to the above expression and deduce the equivalence principle, i.e. gravity equally couples to all the matter. Thus, $\kappa_{1}=...=\kappa_{n}=\kappa$ [^4] This $\kappa$ should not be confused with $\kappa$ used in section (4.1). Hereafter, $\kappa=\sqrt{8\pi G}$.
$\therefore S^{(0,-1)}=\sum_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}.$ (63)
Let us now evaluate the first $\mathcal{O}(l^{-1})$ term in eq. (58). We can not simply perform the $z^{\prime}$-integral to generate the delta function, because the momentum derivative operator is acting on the exponential. Therefore, we first do the integration by parts in $p$ to move the exponential outside the derivative operator, and then perform the $z^{\prime}$-integral as follows:
$\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}\left(-3i\dfrac{p_{i}^{\mu}}{2ml}\dfrac{\partial}{\partial p^{\mu}}\right)e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\\ =\dfrac{3i}{2ml}\int d^{4}z\int\dfrac{d^{4}p}{(2\pi)^{4}}\ p_{i}^{\mu}\dfrac{\partial}{\partial p^{\mu}}\left(\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{-ip\cdot z}\right)\int d^{4}z^{\prime}e^{i(p-p_{i}-k)\cdot z^{\prime}}\\ =\dfrac{3i}{2ml}\int d^{4}z\int\dfrac{d^{4}p}{(2\pi)^{4}}\ p_{i}^{\mu}\dfrac{\partial}{\partial p^{\mu}}\left(\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{-ip\cdot z}\right)(2\pi)^{4}\delta^{(4)}(p-p_{i}-k)\\ =\dfrac{3i}{2ml}\int d^{4}z\left(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2p_{i}\cdot k}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{2}}-i\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(p_{i}\cdot z)}{2p_{i}\cdot k}\right)e^{-ip_{i}\cdot z},$ (64)
where in the last line, the $p$-integral is performed and the soft limit is applied. Similarly, we evaluate the last two $\mathcal{O}(l^{-1})$ terms in eq. (58). The only difference being that we have to do integration by parts twice to move the exponential outside the derivative operators. After evaluating and collecting all the terms, we obtain
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa\int d^{4}z\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{m}{l}\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{2m}{l}\dfrac{p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\bigg{)}e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}(\{p_{i}\})\\ =\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{im}{l}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}+\dfrac{2mi}{l}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg{)}\Gamma^{(0)}_{n}(\{p_{i}\})+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}(\{p_{i}\}).$ (65)
Where,
@@
$A^{(1)}=-im\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg{)}.$ (66)
Thus, the leading correction to the flat space soft factor is
@@
$S^{(1,0)}=-im\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}\partial_{p_{i}}^{\beta}\bigg{)}.$ (67)
Consistency checks:
  • The gauge transformation does not receive any contributions of $\mathcal{O}(l^{-1})$. Performing the gauge transformation, $\varepsilon_{\alpha\beta}\to\varepsilon_{\alpha\beta}-i\xi_{\alpha}k_{\beta}-i\xi_{\beta}k_{\alpha}$, we have verified that the leading correction $S^{(1,0)}$ to the flat space soft factor is gauge invariant.
  • To check the universal feature of the leading correction term, we have used a different set of scalar field modes (13). The new set of modes has no terms of $\mathcal{O}(l^{-1})$, so there are no leading corrections to the flat space soft factor for this set of modes. Since the leading corrections to the soft factor change for different set of modes, it is non-universal.

### v3

@@ 6 Perturbative corrections to flat space soft graviton theorem
We must note that the perturbative corrections to the flat space soft factor are only valid for $\omega l>>1$. And the soft limit of scattering amplitudes in flat space requires $\omega<<p_{i}^{\mu}$, where $p_{i}^{\mu}$ is the momenta of the hard particles. Thus, our analysis, and hence results are valid in the regime $\frac{1}{l}<<\omega<<p_{i}^{\mu}$. For the soft graviton scattering in the de Sitter background, we take the double scaling limit Banerjee_Bhattacharjee_Mitra in which $\omega\to 0$ and $l\to\infty$ simultaneously, keeping their product $\omega l$ finite. In these limits, it is convenient to introduce a new parameter $\delta=1/\omega l$. To put the perturbative corrections in a proper hierarchy of leading, subleading, and sub-subleading corrections, the full soft factor (i.e. flat space + perturbative corrections) can be expanded in $\delta$ as follows:
[…]
where ‘$s$’ represents the power of $\omega$, i.e., $S^{(r,s)}\sim\mathcal{O}(\omega^{s})$. In this soft expansion, $S^{(0,-1)}$ is the flat space leading soft factor, and the higher order terms are leading, subleading, sub-subleading corrections, and so on.
@@ 6.1 Leading corrections
Let us substitute the scalar field propagator (22), scalar field modes (10), and graviton modes (39) into Eq. (61) and simplify the expression by keeping the terms up to $\mathcal{O}(l^{-1})$.
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg[\dfrac{p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}+\dfrac{3i(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2lm(p^{2}+m^{2})}+\dfrac{m(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{l(p^{2}+m^{2})}\\ +\dfrac{5(p_{i}\cdot z^{\prime})(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2lm(p^{2}+m^{2})}-\dfrac{i(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2lm(p^{2}+m^{2})}-\dfrac{imz^{\prime 2}(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2l(p^{2}+m^{2})}\bigg]e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (64)
To simplify the above expression, we replace all $z^{\prime\mu}$ by $i\frac{\partial}{\partial p_{\mu}}$.
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg[\dfrac{p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}+\dfrac{3i(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2lm(p^{2}+m^{2})}-\dfrac{im(p^{\alpha}\varepsilon_{\alpha\beta})}{l(p^{2}+m^{2})}\dfrac{\partial}{\partial p_{\beta}}\\ -\dfrac{5i(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2lm(p^{2}+m^{2})}p_{i}^{\gamma}\dfrac{\partial}{\partial p^{\gamma}}+\dfrac{i(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2lm(p^{2}+m^{2})}p_{i}^{\gamma}p_{i}^{\delta}\dfrac{\partial}{\partial p^{\gamma}}\dfrac{\partial}{\partial p^{\delta}}+\dfrac{im(p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2l(p^{2}+m^{2})}\dfrac{\partial}{\partial p^{\gamma}}\dfrac{\partial}{\partial p_{\gamma}}\bigg]e^{i(p-p_{i}-k)\cdot z^{\prime}}\\ \cdot e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (65)
The above expression contains $\mathcal{O}(l^{0})$ and $\mathcal{O}(l^{-1})$ terms. We analyze them step by step. Let us first evaluate the $\mathcal{O}(l^{0})$ term, which should produce the flat space soft factor.
@@
$\Gamma^{(0)}_{n+1}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{p^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (66)
This expression is trivial to simplify. By performing the $z^{\prime}$-integral in the above expression, one can generate the delta function $\delta^{(4)}(p-p_{i}-k)$, which can be integrated w.r.t. $p$ to replace all $p$ by $p_{i}+k$ in the resultant expression. Finally, by taking the soft limit, one can obtain
@@
$\Gamma^{(0)}_{n+1}(\{p_{1},...,p_{n}\},k)=$ $-i\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\int d^{4}ze^{-ip_{i}\cdot z}$
$\cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle.$ (67)
The momentum $p_{i}$ in the exponential is on-shell, and hence it will remain the same for all the scalars. Thus, we can move $z$-integral outside the summation and rewrite the above expression as
$\Gamma^{(0)}_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\Gamma^{(0)}_{n}{(\{p_{1},...,p_{n}\})},$ (68)
which is the flat space soft graviton theorem. One can apply the Ward identity to the above expression and deduce the equivalence principle, i.e., gravity equally couples to all the matter. Thus, $\kappa_{1}=...=\kappa_{n}=\kappa$.
Therefore, the flat space soft factor is
$A^{(0)}=\sum_{i=1}^{n}\frac{\kappa}{2}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k},$ (69)
and from (63), we can write
$S^{(0,-1)}=\sum_{i=1}^{n}\frac{\kappa}{2}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{\omega p_{i}\cdot\hat{k}}$. (70)
Let us now evaluate the $\mathcal{O}(l^{-1})$ terms in (65). The second term in (65) is trivial like the first term and can be evaluated similarly. However, the third and all the subsequent terms are non-trivial and require integration by parts in $p$ before performing the $z^{\prime}$-integral. We can not simply perform the $z^{\prime}$-integral to generate the delta function, because the momentum derivative operator is acting on the exponential. Therefore, we first do the integration by parts in $p$ to move the exponential outside the derivative operator, and then perform the $z^{\prime}$-integral. For instance, the third term in (65) is evaluated as follows:
$\int d^{4}zd^{4}z^{\prime}$ $\int\dfrac{d^{4}p}{(2\pi)^{4}}\dfrac{p^{\alpha}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{-ip\cdot z}\left(-\dfrac{im}{l}\dfrac{\partial}{\partial p_{\beta}}\right)e^{i(p-p_{i}-k)\cdot z^{\prime}}$
$=\dfrac{im}{l}\int d^{4}z\int\dfrac{d^{4}p}{(2\pi)^{4}}\ \dfrac{\partial}{\partial p_{\beta}}\left(\dfrac{p^{\alpha}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{-ip\cdot z}\right)\int d^{4}z^{\prime}e^{i(p-p_{i}-k)\cdot z^{\prime}}$
$=\dfrac{im}{l}\int d^{4}z\int\dfrac{d^{4}p}{(2\pi)^{4}}\ \dfrac{\partial}{\partial p_{\beta}}\left(\dfrac{p^{\alpha}\varepsilon_{\alpha\beta}}{p^{2}+m^{2}}e^{-ip\cdot z}\right)(2\pi)^{4}\delta^{(4)}(p-p_{i}-k)$
$=\dfrac{im}{l}\int d^{4}z\left(-i\dfrac{p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{2p_{i}\cdot k}-\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{2}}\right)e^{-ip_{i}\cdot z},$ (71)
where in the last line, the $p$-integral is performed and the soft limit is applied. We similarly evaluate all the remaining terms in (65). The only difference is that for the last two terms, we have to do integration by parts twice to put the exponential factor outside the derivative operators. After evaluating and collecting all the terms in (65), we obtain
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=-i\sum_{i=1}^{n}\frac{\kappa}{2}\int d^{4}z\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{m}{l}\dfrac{(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{2m}{l}\dfrac{p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\bigg)e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}(\{p_{1},...,p_{n}\})\\ =\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}-\dfrac{im}{l}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}+\dfrac{2mi}{l}\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg)\Gamma^{(0)}_{n}(\{p_{1},...,p_{n}\})+\dfrac{1}{l}A^{(0)}\Gamma^{(1)}_{n}(\{p_{1},...,p_{n}\}).$ (72)
Where, in the third line, all $z$ coordinates are replaced by $i\partial_{p_{i}}$ to transform the soft factor in the momentum space. The $\mathcal{O}(l^{-1})$ correction to the flat space soft factor is
$A^{(1)}=-im\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}k\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot k}\partial_{p_{i}}^{\beta}\bigg),$ (73)
and from (63), the leading correction to the flat space soft factor is
$S^{(1,0)}=-im\sum_{i=1}^{n}\frac{\kappa}{2}\bigg(\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}\partial_{p_{i}}^{\beta}\bigg).$ (74)
In subsection (6.3), we will see that the leading corrections receive contributions from the diagram where soft graviton is attached to an internal line, i.e., Fig. (3).